Suppose f(x)= 4x-10.Evaluate each of the following.
1. f(2x) 2. f(n) 3. f(-3y) 4. f(t+1) 5. f(t-1)

6. f(r+2) 7. f(c+10)

Answer by sinadone(8)

You just substitute.
For example: f(2x)=4(2x)-10
4 * 2x = 8x-10
Set it equal to zero and solve for x.
8x-10=0
Move the 10 to the right, making it negative.
8x=-10
Then divide by 8.
x=-5/4
You do the same with the rest of the problems.